A rectangle has a diagonal length of 7 cm and a width of 3 cm. What's the area (in simplified radical form).
Answer by jorel1380(3719)

7^2-3^2=L^2
49-9=L^2
L=√40=√4√10=2√10
Area=3x2√10=6√10..